Calculate the mass of the of an object that has d
Answer #1

16. DeBroglie equation,

m = h/lv

where,

m = mass

h = 6.626 x 10^-34 J.s

l = 7.21 x 10^-39 m (wavelength)

v = 163.2 km/hr = 163200/3600 m/s

Feed values,

m = 6.626 x 10^-34/(7.21 x 10^-39 x 45.33) = 2027.3 kg

17.

a. mass equivalent to 6.43 moles of Ca3(PO4)2 = 6.43 x 310.17 = 1994.40 g

b. moles of CO2 = g/molar mass = 63.2/44.01 = 1.44 moles

18. find moles

C = 12.01/12 = 1

H = 4.04/1.01 = 4

Divide by smallest factor

C = 1/1 = 1

H = 4/1 = 4

empirica formula = CH4

mass = 12 + 4 = 16 g

So to get, 23.1 g carbon, we would need = 2.0g of this compound.

19. energy = R[1/ni^2-1/nf^2]

with,

ni = 6

nf = 3

we have,

energy = 2.18 x 10^-18[1/36-1/9]

= -1.82 J